[Hallc_running] [New Logentry] RC Report

smithg at jlab.org smithg at jlab.org
Tue Nov 19 09:35:02 EST 2019

Logentry Text:
RC Report for Tuesday, Nov. 19, 2019

>From today’s 8 AM meeting, and input from JP:

The work to swap out the leaky 1L10 girder in the North Linac (NL) was completed yesterday. Cryomodule 8 is stung and at 4K. The other 2 CMs (9 & 10, at either end of the 1L10 girder) will be stung and cooled tomorrow pending a successful 1L10 leak check today. Bottom line: the MCC schedule is on track, and the same as described yesterday: beam to C on the 2nd.

I caught JP on his way to the safety meeting. He reports that all the fiber testing is done! Woo-hoo! The one bad fiber mentioned yesterday is at the vendor for repair and is supposed to be back here today, at which point we'll have all 4+4+2 fibers. His plan is to install them all at once Wed/Thu so the LOSP inspection can happen on Friday this week 11/22. Then this weekend and next week the plan is to deal with any findings from this inspection, get the LOSP approved, and get on with the alignment. He also reported that the instabilities in the target motion have been fixed. All in all very good news. Bottom line: he's still optimistic the installation can finish by Dec. 2, with the caveat that target optimization may still require hall access the week of the 2nd. 

To avoid interfering with this tight installation schedule, Thia made the call that Hall C will not participate in the spin dance proposed for 11/29-30. 

The DOE safety team is on site today through the end of the week. Also, the switch to two-factor authentication for physics division happens today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3742197
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20191119/35817350/attachment.html>

More information about the Hallc_running mailing list